How do you graph the equation 6-3/8y=4 ?
Also written as: 6-(3/8)y=4
Where does the graph of the above equation cross the graphed line of (2/3)x + (1/4) y = 4?

Answer by stanbon(75887)

How do you graph the equation 6-3/8y=4 ?
Also written as: 6-(3/8)y=4
------
Solve for "y":
(3/8)y = 2
y = 16/3 = 5 1/3
the graph is a horizontal line thru all points where y = 5 1/3
------------

Where does the graph of the above equation cross the graphed line of (2/3)x + (1/4) y = 4?
Solve for "y":
(1/4)y = (-2/3)x + 4
---
y = (-8/3)x + 16
---
Substitute for "y" and solve for "x":
16/3 = (-8/3)x + 16
-----
16 = -8x + 48
8x = 32
x = 4
-----
Point of intersection: (4,16/3)
====================================
Cheers,
Stan H.